OpenStudy (anonymous):

find the zeros of y=x^2 +9x+8,then use the zeros to find the vertex of the graph of the parabola

OpenStudy (sriram):

x^2 +9x +8 = (x+8) * (x+1) hence zeroes at -1 &-8 hence vertex at -9/2 (mean of the 2 zeroes

OpenStudy (anonymous):

(x+8)(x+1)=0 -8,-1 the x=coordinate of the vertex is in the middle, -4.5 y-coordinate is -12.25 by substituting -4.5 in for x (-4.5,-12.25)